OpenStudy (anonymous):

I have a Differential Equation question... I've just started, so the question is quite simple: How do I know if a function is linear or not? For instance: (x-1)y'' + yxy' - cos x = 1. Is this a linear or non linear function?

OpenStudy (anonymous):

What function?

OpenStudy (anonymous):

In general, a function is linear if it satisfies: f(a+b) = f(a)+f(b) and f(kx) = kf(x) where k is a real number.

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."

OpenStudy (amistre64):

(x-1)y'' + yxy' - cos x = 1 as long as none of the y parts have an exponent greater then 1, or are modified by trig or ln or something unnatural ... it should be good

OpenStudy (amistre64):

y'' is a derivative, of order 2 y' is a derivative of order 1 y is a function that the derivatives came from.

OpenStudy (amistre64):

and in this case, it looks like the y' y part is pushing this to nonlinear